find the annual interest rate to the nearest tenth of a percent for 1100 to grow to 1800 if the interest is compounded quarterly for 6 years
Answer by stanbon(75887)

find the annual interest rate to the nearest tenth of a percent for 1100 to grow to 1800 if the interest is compounded quarterly for 6 years
------
A(t) = P(1+(r/n))^(nt)
------
1800 = 1100(1+(r/4))^(4*6)
-----
(1+(r/4)^24 = 18/11
----
24*log(1+(r/4)) = log(18)-log(11)
-----------------
log(1+(r/4)) = 0.0089
----
1+(r/4) = 10*0.0089
---
1 + (r/4) = 1.0207
---
r/4 = 0.0207
---
r = 0.0829
================
Cheers,
Stan H.
=======================